OpenStudy (ebonyserrano):

A total of 408 tickets were sold for the school play. They were either adult tickets or student tickets. The number of student tickets sold was three times the number of adult tickets sold. How many adult tickets were sold?

OpenStudy (ebonyserrano):

@Will.H

OpenStudy (will.h):

Adult tickets = x Students tickets= 3x Total = x + 3x = 408 4x = 408 X = 102

OpenStudy (shadowlegendx):

Do you understand how Will got that?

OpenStudy (shadowlegendx):

Let x = amount of adult tickets sold Let y = amount of student tickets sold Since there are only these two types of tickets, we can infer that x + y = 408 Using this bit of information: "The number of student tickets sold was three times the number of adult tickets sold." We can infer that, y = 3x So going back to our first equation, we can replace y with 3x x + 3x = 408 Like terms 4x = 408 Divide the equation by 4 x = 102 102 adult tickets were sold